Let x be Bin(n,p). Find the mean square error of the p parameter estimator




1
Expert's answer
2022-01-17T16:27:43-0500

"MSE(\hat p)=E[(\hat p-p)^2]=E[(X/n-p)^2]=E[X^2/n^2-2pX/n+p^2]="


"=E[X^2]\/n^2-2pE[X]\/n+p^2=np((n-1)p+1)\/n^2-2np^2\/n+p^2="


"=p(1-p)\/n"